# Implicit Bias of Per-sample Adam on Separable Data: Departure from the Full-batch Regime

### v2

@@ 4 Generalization: AdamProxy
describes the implicit bias of the steepest descent algorithm with respect to the $\ell_{p}$-norm in linear classification tasks (Gunasekar et al., 2018a). However, Equation 5 does not correspond to the KKT conditions of a conventional optimization problem. To address this, we introduce a novel framework to describe the convergence direction, based on a parametric optimization problem combined with fixed-point analysis between dual variables.
@@
Theorem 4.8 shows how the parametric optimization problem $P_{\text{Adam}}({\mathbf{c}})$ captures the characterization from Lemma 4.5. The central idea is to treat the vector $\mathbf{c}$} from Equation 5 in a dual role: as both the parameter of $P_{\text{Adam}}({\mathbf{c}})$ and as its corresponding dual variable. The convergence direction is then identified at the point where these two roles coincide, leading naturally to the fixed-point formulation.
@@
Example 4.9 (Revisiting GR data).
For GR data $\{{\mathbf{x}}_{i}\}_{i\in[N]}$, the matrix ${\mathbf{M}}({\mathbf{c}})$ reduces to a scaled identity for every ${\mathbf{c}}\in\Delta^{N-1}$. Hence, the parametric optimization problem $P_{\text{Adam}}({\mathbf{c}})$ narrows down to the standard SVM formulation
@@
Therefore, Theorem 4.8 implies that AdamProxy converges to the $\ell_{2}$-max-margin solution. This finding is consistent with Theorem 3.3, which establishes the directional convergence of Inc-Adam on GR data. Together, these results indicate that the structural property of GR data that eliminates coordinate adaptivity persists in the limit $\beta_{2}\to 1$.
@@
We next validate the fixed-point characterization in Theorem 4.8 using the Gaussian dataset from Figure 1. The theoretical limit direction is given by the fixed point of $T$ defined in Theorem 4.8, which we compute via the iteration in Algorithm 3. As shown in Figure 3, both AdamProxy and mini-batch Adam variants with batch size $1$ converge to the predicted solution, confirming the fixed-point formulation and the effectiveness of Algorithm 3. Furthermore, this demonstrates that, depending on the dataset, the limit direction of mini-batch Adam may differ from both the conventional $\ell_{2}$- and $\ell_{\infty}$-max-margin solutions.
Figure 3: Mini-batch Adam converges to the fixed-point solution on Gaussian data. We train on the same Gaussian data as in Figure 1 and plot the cosine similarity of the weight vector with the $\ell_{2}$-max-margin solution (left) and the fixed-point solution (right). The results show that variants of mini-batch Adam with batch size $1$ converge to the fixed-point solution obtained by Algorithm 3, consistent with our theoretical prediction (Theorem 4.8).
@@
has the solution $\hat{{\mathbf{w}}}_{\infty}=(\frac{1}{x_{0}+(d-1)\delta},\cdots,\frac{1}{x_{0}+(d-1)\delta})\in\mathbb{R}^{d}$. Notice that ${\mathbf{c}}^{*}=(1,0,\cdots,0)\in\Delta^{d-1}$ is a fixed point of $T$ in Theorem 4.8, and $\hat{{\mathbf{w}}}_{\infty}={\mathbf{p}}({\mathbf{c}}^{*})$; detailed calculations are deferred to Appendix F. Consequently, the $\ell_{\infty}$-max-margin solution serves a candidate for the convergence direction of AdamProxy as predicted by Theorem 4.8. To verify this, we run AdamProxy and mini-batch Adam variants with batch size $1$ on shifted-diagonal data given by ${\mathbf{x}}_{0}=(1,\delta,\delta,\delta)$, ${\mathbf{x}}_{1}=(\delta,2,\delta,\delta)$, ${\mathbf{x}}_{2}=(\delta,\delta,4,\delta)$, and ${\mathbf{x}}_{3}=(\delta,\delta,\delta,8)$ with $\delta=0.1$. As shown in Figure 4, all mini-batch Adam variants converge to the $\ell_{\infty}$-max-margin solution, consistent with the theoretical prediction.
Figure 4: Mini-batch Adam converges to the $\ell_{\infty}$-max-margin solution on a shifted-diagonal dataset. We train on the dataset ${\mathbf{x}}_{0}=(1,\delta,\delta,\delta)$, ${\mathbf{x}}_{1}=(\delta,2,\delta,\delta)$, ${\mathbf{x}}_{2}=(\delta,\delta,4,\delta)$, and ${\mathbf{x}}_{3}=(\delta,\delta,\delta,8)$ with $\delta=0.1$. Variants of mini-batch Adam with batch size $1$ converge to the $\ell_{\infty}$-max-margin direction.
A key limitation of our analysis is that it assumes $\beta_{2}\to 1$ and a batch size of $1$. In Appendix A, we provide a preliminary analysis of how batch size and momentum hyperparameters affect the implicit bias of mini-batch Adam. In particular, Section A.2 explains why our fixed-point framework does not directly extend to finite $\beta_{2}$.
@@ 5 Signum can Retain $\ell_{\infty}$-bias under Mini-batch Regime
and such $\epsilon$ is given by
$\epsilon=\begin{cases}\frac{1}{2D\cdot\tfrac{N}{b}(\tfrac{N}{b}-1)}\min\!\left\{\delta,\tfrac{\gamma_{\infty}}{2}\right\}&\text{if }b<N,\\ 1&\text{if }b=N.\end{cases}$
@@ 6 Related Work
Adam (Kingma and Ba, 2015) and its variant AdamW (Loshchilov and Hutter, 2019) are standard optimizers for large-scale models, particularly in domains like language modeling where SGD often falls short. A significant body of research seeks to explain this empirical success. One line focuses on convergence guarantees. The influential work of Reddi et al. (2018) demonstrates Adam’s failure to converge on certain convex problems, which motivates numerous studies establishing its convergence under various practical conditions (Défossez et al., 2022; Zhang et al., 2022; Li et al., 2023; Hong and Lin, 2024; Ahn and Cutkosky, 2024; Jin et al., 2025). Another line investigates why Adam outperforms SGD, attributing its success to robustness against heavy-tailed gradient noise (Zhang et al., 2020), better adaptation to ill-conditioned landscapes (Jiang et al., 2023; Pan and Li, 2023), and effectiveness in contexts of heavy-tailed class imbalance or gradient/Hessian heterogeneity (Kunstner et al., 2024; Zhang et al., 2024b; Tomihari and Sato, 2025). Ahn et al. (2024) further observe that this performance gap arises even in shallow linear Transformers.
@@
Recent work increasingly examines Adam’s implicit bias and its connection to $\ell_{\infty}$-geometry. This link is motivated by Adam’s similarity to SignGD (Balles and Hennig, 2018; Bernstein et al., 2018), which performs normalized steepest descent under the $\ell_{\infty}$-norm. Kunstner et al. (2023) show that the performance gap between Adam and SGD increases with batch size, while SignGD achieves performance similar to Adam in the full-batch regime, supporting this connection. Zhang et al. (2024a) prove that Adam without a stability constant converges to the $\ell_{\infty}$-max-margin solution in separable linear classification, later extended to multi-class classification by Fan et al. (2025). Complementing these results, Xie and Li (2024) show that AdamW implicitly solves an $\ell_{\infty}$-norm-constrained optimization problem, connecting its dynamics to the Frank-Wolfe algorithm. Exploiting this $\ell_{\infty}$-geometry is argued to be a key factor in Adam’s advantage over SGD, particularly for language model training (Xie et al., 2025).
@@ 7 Discussion and Future Work
Empirical evidence shows that Adam’s advantage over SGD is most pronounced in large-batch training, while the gap diminishes with smaller batches (Kunstner et al., 2023; Srećković et al., 2025). Our results suggest a possible explanation: the $\ell_{\infty}$-adaptivity of Adam, proposed as the source of its advantage (Xie et al., 2025), may vanish in the mini-batch regime. An important direction for future work is to investigate whether this loss of $\ell_{\infty}$-adaptivity extends beyond linear models and how it interacts with practical large-scale training.
@@
Our analysis for general dataset relies on the asymptotic regime $\beta_{2}\to 1$ and on incremental Adam as a tractable surrogate. Extending the framework to finite $\beta_{2}$, larger batch sizes, and common sampling schemes (e.g., random reshuffling) would make the theory more complete. See Appendix A for further discussion. Relaxing technical assumptions and developing tools that apply under broader conditions also remain important directions.
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grants funded by the Korean government (MSIT) (No. RS-2023-00211352; No. RS-2024-00421203).
@@ A.1 Effect of Hyperparameters on Mini-batch Adam
Figure 7: $\beta_{1}$ does not affect the convergence direction of mini-batch Adam for large $\beta_{2}$. We train on the same Gaussian data as in Figure 1, varying $\beta_{1}\in\{0.9,0.5,0.1\}$ with fixed $\beta_{2}=0.95$, and plot the cosine similarity between the weight vector and the fixed-point solution (Algorithm 3). All mini-batch Adam variants with batch size 11 consistently converge to the fixed-point solution.
Figure 8: $\beta_{2}$ affects the convergence direction of mini-batch Adam. We train on the same Gaussian data as in Figure 1, varying $\beta_{2}\in\{0.9,0.5,0.1\}$ with fixed $\beta_{1}=0.1$, and plot the cosine similarity between the weight vector and the fixed-point solution (Algorithm 3). Mini-batch Adam variants with batch size 11 deviate increasingly from the fixed-point solution as $\beta_{2}$ decreases.
@@ Appendix B Additional Experiments
To investigate the universality of Theorem 3.3 with respect to the choice of the momentum hyperparameters, we run mini-batch Adam (with batch size 1) on GR dataset ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{3}=(4,-4,4,-4)$, varying the momentum hyperparameters $(\beta_{1},\beta_{2})\in\{(0.1,0.1),(0.5,0.5),(0.9,0.95)\}$. Figure 9 demonstrates that its limiting behavior toward $\ell_{2}$-max-margin solution consistently holds on the broad choices of $(\beta_{1},\beta_{2})$.
@@
Figure 9: Mini-batch Adam converges to the max $\ell_{2}$-margin solution for GR data. We train on GR dataset ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{3}=(4,-4,4,-4)$, varying the momentum hyperparameters. In all tested configurations, the family of mini-batch Adam algorithms with batch size $1$ converge to the $\ell_{2}$ max-margin solution, which deviate significantly from the $\ell_{\infty}$ bias of full-batch Adam.
@@ Appendix C Experimental Details
  • Generalized Rademacher (GR) data (Figures 2 and 9): We use ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{3}=(4,-4,4,-4)$.
@@
We minimize the exponential loss using various algorithms. Momentum hyperparameters are $(\beta_{1},\beta_{2})=(0.9,0.95)$ for Adam and $\beta=0.99$ for Signum unless specified otherwise. For Adam and Signum variants, we use a learning rate schedule $\eta_{t}=\eta_{0}(t+2)^{-a}$ with $\eta_{0}=0.1$ and $a=0.8$, following our theoretical analysis. Gradient descent uses a fixed learning rate $\eta_{t}=\eta_{0}=0.1$. Margins with respect to different norms are computed using CVXPY [Diamond and Boyd, 2016].
@@ Appendix D Missing Proofs in Section 2
In this section, we provide the omitted proofs in Section 2, which describes asymptotic behaviors of Det-Adam and Inc-Adam. We first introduce Lemma D.1 originated from Zou et al. [2023, Lemma A.2], which gives a coordinate-wise upper bound of updates of both Det-Adam and Inc-Adam. Then, we prove Propositions 2.4 and 2.5 by approximating two momentum terms.
@@
Following the proof of Zou et al. [2023, Lemma A.2], we can easily show that the given upper bound holds for both Det-Adam and Inc-Adam. We prove the case of Inc-Adam, while it naturally extends to Det-Adam. By Cauchy-Schwartz inequality, we get
@@
We recall Lemma 6.1 in Zhang et al. [2024a], stating that

### v3

@@ 4 Generalization: AdamProxy
describes the implicit bias of the steepest descent algorithm with respect to the $\ell_{p}$-norm in linear classification tasks (Gunasekar et al., 2018b). However, Equation 5 does not correspond to the KKT conditions of a conventional optimization problem. To address this, we introduce a novel framework to describe the convergence direction, based on a parametric optimization problem combined with fixed-point analysis between dual variables.
@@
Theorem 4.8 shows how the parametric optimization problem $P_{\text{Adam}}({\mathbf{c}})$ captures the characterization from Lemma 4.5. The central idea is to treat the vector $\mathbf{c}$} from Equation 5 in a dual role: as both the parameter of $P_{\text{Adam}}({\mathbf{c}})$ and as its corresponding dual variable. The convergence direction is then identified at the point where these two roles coincide, leading naturally to the fixed-point formulation. Detailed proofs are deferred to Appendix F.
@@
Example 4.9 (Revisiting SR data).
For SR data $\{{\mathbf{x}}_{i}\}_{i\in[N]}$, the matrix ${\mathbf{M}}({\mathbf{c}})$ reduces to a scaled identity for every ${\mathbf{c}}\in\Delta^{N-1}$. Hence, the parametric optimization problem $P_{\text{Adam}}({\mathbf{c}})$ narrows down to the standard SVM formulation
@@
Therefore, Theorem 4.8 implies that AdamProxy converges to the $\ell_{2}$-max-margin solution. This finding is consistent with Theorem 3.3, which establishes the directional convergence of Inc-Adam on SR data. Together, these results indicate that the structural property of SR data that eliminates coordinate adaptivity persists in the limit $\beta_{2}\to 1$.
@@
We next validate the fixed-point characterization in Theorem 4.8 using the Gaussian dataset from Figure 1. The theoretical limit direction is given by the fixed point of $T$ defined in Theorem 4.8, which we compute via the iteration in Algorithm 3. As shown in Figure 3, both AdamProxy and mini-batch Adam variants with a batch size of $1$ converge to the predicted solution, confirming the fixed-point formulation and the effectiveness of Algorithm 3. Furthermore, this demonstrates that, depending on the dataset, the limit direction of mini-batch Adam may differ from both the conventional $\ell_{2}$- and $\ell_{\infty}$-max-margin solutions.
Figure 3: Mini-batch Adam converges to the fixed-point solution on Gaussian data. We train on the same Gaussian data as in Figure 1 and plot the cosine similarity of the weight vector with the $\ell_{2}$-max-margin solution (left) and the fixed-point solution (right). The results show that variants of mini-batch Adam with a batch size of $1$ converge to the fixed-point solution obtained by Algorithm 3, consistent with our theoretical prediction (Theorem 4.8).
@@
has the solution $\hat{{\mathbf{w}}}_{\infty}=(\frac{1}{x_{0}+(d-1)\delta},\cdots,\frac{1}{x_{0}+(d-1)\delta})\in\mathbb{R}^{d}$. Notice that ${\mathbf{c}}^{*}=(1,0,\cdots,0)\in\Delta^{d-1}$ is a fixed point of $T$ in Theorem 4.8, and $\hat{{\mathbf{w}}}_{\infty}={\mathbf{p}}({\mathbf{c}}^{*})$; detailed calculations are deferred to Appendix F. Consequently, the $\ell_{\infty}$-max-margin solution serves a candidate for the convergence direction of AdamProxy as predicted by Theorem 4.8. To verify this, we run AdamProxy and mini-batch Adam variants with a batch size of $1$ on shifted-diagonal data given by ${\mathbf{x}}_{0}=(1,\delta,\delta,\delta)$, ${\mathbf{x}}_{1}=(\delta,2,\delta,\delta)$, ${\mathbf{x}}_{2}=(\delta,\delta,4,\delta)$, and ${\mathbf{x}}_{3}=(\delta,\delta,\delta,8)$ with $\delta=0.1$. As shown in Figure 4, all mini-batch Adam variants converge to the $\ell_{\infty}$-max-margin solution, consistent with the theoretical prediction.
Figure 4: Mini-batch Adam converges to the $\ell_{\infty}$-max-margin solution on a shifted-diagonal dataset. We train on the dataset ${\mathbf{x}}_{0}=(1,\delta,\delta,\delta)$, ${\mathbf{x}}_{1}=(\delta,2,\delta,\delta)$, ${\mathbf{x}}_{2}=(\delta,\delta,4,\delta)$, and ${\mathbf{x}}_{3}=(\delta,\delta,\delta,8)$ with $\delta=0.1$. Variants of mini-batch Adam with a batch size of $1$ converge to the $\ell_{\infty}$-max-margin direction.
A key limitation of our analysis is that it assumes $\beta_{2}\to 1$ and a batch size of $1$. In Appendix A, we provide a preliminary analysis of how batch size and momentum hyperparameters affect the implicit bias of mini-batch Adam. In particular, Section A.2 explains why our fixed-point framework does not directly extend to general $\beta_{2}<1$.
@@ 5 Signum can Retain $\ell_{\infty}$-bias under Mini-batch Regime
and
$\epsilon=\frac{1}{2D\cdot\tfrac{N}{b}(\tfrac{N}{b}-1)}\min\!\left\{\delta,\tfrac{\gamma_{\infty}}{2}\right\}\quad\text{if }b<N,\qquad\epsilon=1\quad\text{if }b=N.$
@@ 6 Related Work
Adam (Kingma and Ba, 2015) and its variant AdamW (Loshchilov and Hutter, 2019) are standard optimizers for large-scale models, particularly in domains like language modeling where SGD often falls short. A significant body of research seeks to explain this empirical success. One line focuses on convergence guarantees. The influential work of Reddi et al. (2018) demonstrates Adam’s failure to converge on certain convex problems, which motivates numerous studies establishing its convergence under various practical conditions (Défossez et al., 2022; Zhang et al., 2022; Li et al., 2023; Hong and Lin, 2024; Ahn and Cutkosky, 2024; Jin et al., 2025). Another line investigates why Adam outperforms SGD, attributing its success to robustness against heavy-tailed gradient noise (Zhang et al., 2020), better adaptation to ill-conditioned landscapes (Jiang et al., 2023; Pan and Li, 2023), and effectiveness in contexts of heavy-tailed class imbalance or gradient/Hessian heterogeneity (Kunstner et al., 2024; Zhang et al., 2024b; Tomihari and Sato, 2025). Ahn et al. (2024) further observe that this performance gap arises even in shallow linear Transformers. Recent works investigate how the choice of momentum hyperparameters (Orvieto and Gower, 2025) and the rotation operation (Zhang et al., 2025) affect the performance of Adam.
@@
A growing body of work examines Adam’s implicit bias and its connection to $\ell_{\infty}$-geometry. This link is motivated by Adam’s similarity to SignGD (Balles and Hennig, 2018; Bernstein et al., 2018), which performs normalized steepest descent under the $\ell_{\infty}$-norm. Kunstner et al. (2023) show that the performance gap between Adam and SGD increases with batch size, while SignGD achieves performance similar to Adam in the full-batch regime, supporting this connection. Zhang et al. (2024a) prove that Adam without a stability constant converges to the $\ell_{\infty}$-max-margin solution in separable linear classification, later extended to multi-class classification by Fan et al. (2025). Tsilivis et al. (2025) investigate implicit bias of steepest descent in homogeneous neural networks, supporting that SignGD describes a typical dynamics of Adam. Complementing these results, Xie and Li (2024) show that AdamW implicitly solves an $\ell_{\infty}$-norm-constrained optimization problem, connecting its dynamics to the Frank-Wolfe algorithm. Exploiting this $\ell_{\infty}$-geometry is argued to be a key factor in Adam’s advantage over SGD, particularly for language model training (Xie et al., 2025). Vasudeva et al. (2025) examine how Adam and GD show different implicit biases when training two-layer ReLU networks, describing Adam’s richer and more diverse decision boundary.
@@ 7 Discussion and Future Work
Empirical evidence shows that Adam’s advantage over SGD is most pronounced in large-batch training, while the gap diminishes with smaller batches (Kunstner et al., 2023; Srećković et al., 2025; Marek et al., 2025). Our results suggest a possible explanation: the $\ell_{\infty}$-adaptivity of Adam, proposed as the source of its advantage (Xie et al., 2025), may vanish in the mini-batch regime. An important direction for future work is to investigate whether this loss of $\ell_{\infty}$-adaptivity extends beyond linear models and how it interacts with practical large-scale training.
@@
Our analysis for general dataset relies on the asymptotic regime $\beta_{2}\to 1$ and on incremental Adam as a tractable surrogate. Extending the framework to incorporate general $\beta_{2}<1$, larger batch sizes, and common sampling schemes (e.g., random reshuffling) would make the theory more complete. See Appendix A for further discussion. Developing additional theoretical tools that can be applied under weaker assumptions also remains an important direction.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grants funded by the Korean government (MSIT) (No. RS-2023-00211352; No. RS-2024-00421203) and the InnoCORE program of the Ministry of Science and ICT (No. N10250156).
Reproducibility Statement
All assumptions and theorems for our theoretical results are stated in the main paper, with their complete proofs deferred to Appendices D, E, F and G. The primary experimental setups are described in the main paper and Appendix C, while details for supplementary experiments are provided in Appendices B and C.
Declaration of LLM Usage
The authors utilized LLMs to improve the grammar and readability of this manuscript. The core conceptualization, analysis, and writing of the content were performed exclusively by the authors.
@@ A.1 Effect of Hyperparameters on Mini-batch Adam
Figure 7: $\beta_{1}$ does not affect the convergence direction of mini-batch Adam for large $\beta_{2}$. We train on the same Gaussian data as in Figure 1, varying $\beta_{1}\in\{0.9,0.5,0.1\}$ with fixed $\beta_{2}=0.95$, and plot the cosine similarity between the weight vector and the fixed-point solution (Algorithm 3). All mini-batch Adam variants with a batch size of 11 consistently converge to the fixed-point solution.
Figure 8: $\beta_{2}$ affects the convergence direction of mini-batch Adam. We train on the same Gaussian data as in Figure 1, varying $\beta_{2}\in\{0.9,0.5,0.1\}$ with fixed $\beta_{1}=0.1$, and plot the cosine similarity between the weight vector and the fixed-point solution (Algorithm 3). Mini-batch Adam variants with a batch size of 11 deviate increasingly from the fixed-point solution as $\beta_{2}$ decreases.
@@ Appendix B Additional Experiments
To investigate the universality of Theorem 3.3 with respect to the choice of the momentum hyperparameters, we run mini-batch Adam (with batch size 1) on SR dataset ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{3}=(4,-4,4,-4)$, varying the momentum hyperparameters $(\beta_{1},\beta_{2})\in\{(0.1,0.1),(0.5,0.5),(0.9,0.95)\}$. Figure 9 demonstrates that its limiting behavior toward $\ell_{2}$-max-margin solution consistently holds on the broad choices of $(\beta_{1},\beta_{2})$.
@@
Figure 9: Mini-batch Adam converges to the max $\ell_{2}$-margin solution for SR data. We train on SR dataset ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{3}=(4,-4,4,-4)$, varying the momentum hyperparameters. In all tested configurations, the family of mini-batch Adam algorithms with a batch size of $1$ converge to the $\ell_{2}$ max-margin solution, which deviate significantly from the $\ell_{\infty}$ bias of full-batch Adam.
@@ Appendix C Experimental Details
  • Scaled Rademacher (SR) data (Figures 2 and 9): We use ${\mathbf{x}}_{0}=(1,1,1,1)$, ${\mathbf{x}}_{1}=(2,2,2,-2)$, ${\mathbf{x}}_{2}=(3,3,-3,-3)$, and ${\mathbf{x}}_{3}=(4,-4,4,-4)$.
@@
We minimize the exponential loss using various algorithms. Momentum hyperparameters are $(\beta_{1},\beta_{2})=(0.9,0.95)$ for Adam and $\beta=0.99$ for Signum unless specified otherwise. For Adam and Signum variants, we use a learning rate schedule $\eta_{t}=\eta_{0}(t+2)^{-a}$ with $\eta_{0}=0.1$ and $a=0.8$, following our theoretical analysis. Gradient descent uses a fixed learning rate $\eta_{t}=\eta_{0}=0.1$. Margins with respect to different norms are computed using CVXPY (Diamond and Boyd, 2016).
@@ Appendix D Missing Proofs in Section 2
In this section, we provide the omitted proofs in Section 2, which describes asymptotic behaviors of Det-Adam and Inc-Adam. We first introduce Lemma D.1 originated from Zou et al. (2023, Lemma A.2), which gives a coordinate-wise upper bound of updates of both Det-Adam and Inc-Adam. Then, we prove Propositions 2.4 and 2.5 by approximating two momentum terms.
@@
Following the proof of Zou et al. (2023, Lemma A.2), we can easily show that the given upper bound holds for both Det-Adam and Inc-Adam. We prove the case of Inc-Adam, while it naturally extends to Det-Adam. By Cauchy-Schwartz inequality, we get
@@
We recall Lemma 6.1 in Zhang et al. (2024a), stating that